### v2

# How Catastrophic is Your LLM? Certifying Risk in Conversation

Abstract
Warning: This paper may contain harmful model outputs.
Large Language Models (LLMs) can produce catastrophic responses in conversational settings that pose serious risks to public safety and security. Existing evaluations often fail to fully reveal these vulnerabilities because they rely on fixed attack prompt sequences, lack statistical guarantees, and do not scale to the vast space of multi-turn conversations. In this work, we propose C3LLM, a novel, principled statistical Certification framework for Catastrophic risks in multi-turn Conversation for LLMs that bounds the probability of an LLM generating catastrophic responses under multi-turn conversation distributions with statistical guarantees. We model multi-turn conversations as probability distributions over query sequences, represented by a Markov process on a query graph whose edges encode semantic similarity to capture realistic conversational flow, and quantify catastrophic risks using confidence intervals. We define several inexpensive and practical distributions—random node, graph path, and adaptive with rejection. Our results demonstrate that these distributions can reveal substantial catastrophic risks in frontier models, with certified lower bounds as high as 70% for the worst model, highlighting the urgent need for improved safety training strategies in frontier LLMs.
1 Introduction
Large Language Models (LLMs) can be used for both beneficial and harmful purposes, ranging from accelerating scientific discovery (Wysocki et al., 2024; Pal et al., 2023) to facilitating the design of bioweapons (Sandbrink, 2023). Although modern LLMs are trained with safety mechanisms (Ouyang et al., 2022; Bai et al., 2022) that are intended to reject unsafe queries, the risk of catastrophic outcomes remains. Catastrophic outcomes refer to highly dangerous or socially damaging responses, such as instructions for building explosives, synthesizing biological weapons, or conducting cyberattacks (SESSION, 2025). While single-turn jailbreak attacks have been widely explored (Yu et al., 2023; Zou et al., 2023; Liu et al., 2024), real-world conversations are inherently multi-turn: an adversary can embed malicious intent in a conversation, gradually steering the model towards harmful content while each query appears innocuous.
Motivation. Most prior works evaluate LLM safety empirically by measuring attack success rates on fixed datasets of query sequences  (Russinovich et al., 2025; Ren et al., 2024). Although informative, these studies have two fundamental limitations: (1) their conclusions depend heavily on the fixed attack sequences, leaving open the possibility that unseen sequences could succeed; and (2) they provide no statistical guarantees, making their findings unreliable and non-generalizable across the vast space of possible conversations. For example, suppose a benchmark provides 20 attack sequences of length 5. In the best case, this benchmark can reveal at most 20 distinct catastrophic behaviors. By contrast, the full conversation space can be much larger: if we construct sequences of length 5 by uniformly combining individual queries from the 20 benchmark sequences, the space contains up to $100^{5}$ possible sequences. Exhaustive testing is infeasible in a large space. Furthermore, not all sequences are equally important; some sequences are more likely than others to trigger catastrophic responses or to represent realistic conversations with a user. Therefore, we want to provide guarantees with respect to probabilistic distributions defined over a large space of multi-turn conversations. Using these guarantees, we can build novel metrics to reliably compare the safety of different frontier models.
Figure 1: Overall statistical certification framework C3LLM(Certification of Catastrophic risks in multi-turn Conversation for LLMs). Starting from a query set, we construct a graph in which edges connect semantically similar queries. On this graph, we define formal specifications as probability distributions over query sequences. For each sampled sequence, we query the LLM, use a judge model to determine whether the response is harmful, and aggregate the results to compute statistical certification bounds on the probability of catastrophic risk.
Challenges. First, existing works on formal guarantees on neural networks typically rely on perturbation analysis within a local neighborhood (e.g., a $l_{\infty}$-ball around the input) (Singh et al., 2025), but such approaches do not naturally apply to prompt-based attacks on LLMs. Second, the catastrophic risk in multi-turn conversations is a temporal property, making it more complex to specify and certify than the single-step settings considered in the literature. Finally, to capture realistic adversarial behavior, we want to define probability distributions that (i) capture realistic conversations that can be exploited by an adversary and (ii) allow distribution shifts, reflecting how real-world attackers adapt their next query based on previous responses from LLMs. Formally specifying and certifying such quantitative, probabilistic, and temporal properties for LLMs has not been attempted before.
This work. When considering a large space, for any LLM, it is possible to find a conversation where the LLM produces catastrophic output. Therefore, qualitative guarantees, i.e., checking whether there exists a single catastrophic conversation, do not lead to a meaningful metric for comparing LLMs. We aim for quantitative guarantees: measuring the probability of catastrophic responses on a randomly sampled conversation. Since exact probabilities cannot be computed in practice (Chaudhary et al., 2024), we focus on high-confidence bounds on this risk through statistical certification.
Benefits of certification over benchmarking. With certification, we bound the probability of catastrophic outputs across all possible sequences with statistical guarantees, not just those in a fixed set of benchmarks. For our previous example, if a certification procedure reports a high-confidence interval of $[0.4,0.6]$ for catastrophic risk, it implies that with high confidence, at least $0.4\times n$ sequences can trigger catastrophic outcomes, where $n$ is the number of samples in the distribution that can be up to $100^{5}$. By reasoning about the entire distribution over queries rather than evaluating only fixed sequences, we can uncover substantially more extensive vulnerabilities.
Main contributions. In this work, we present C3LLM, the first framework (shown in Figure 1) for certifying catastrophic risks in multi-turn conversations with LLMs. We are the first to formally specify the temporal safety of LLM responses in a conversational setting. We provide a general recipe for designing such specifications based on Markov processes on graph representations. We instantiate the framework with three different distributions—random node, graph path, and adaptive with rejection (Section 3), capturing a large number of realistic conversations exploitable by adversaries with fixed or adaptive attack strategies. C3LLM then certifies the target LLM by generating high-confidence bounds on the probability of catastrophic risks for a randomly sampled conversation from the distribution. Our main contributions are:
  • We are the first to design a general recipe for formally specifying the risk of catastrophic responses from LLMs in multi-turn conversations. Conversations are represented as query sequences in a graph where edges encode semantic similarity. We introduce a Markov process over this graph. We instantiate with three representative distributions—random node, graph path, and adaptive with rejection, to reflect both semantic relationships and adaptive attacker behavior.
  • We introduce the first framework for certifying catastrophic risk in multi-turn LLM conversations. We model attacks as probability distributions over query sequences and draw independent and identically distributed (i.i.d.) samples from these distributions. This enables statistical guarantees over vast conversational spaces, providing principled certification of catastrophic risks.
  • We find a non-trivial lower bound on the probability of catastrophic risks across different frontier LLMs. We find that Claude-Sonnet-4 and Nova Premier are the safest while Mistral-Large and DeepSeek-R1 exhibit the highest risks. We conduct case studies to identify common patterns, distractors (additional benign queries in the dialogue making refusals less likely) and context (preceding turns providing supporting information and making harmful targets clearer), that lead to catastrophic outputs.
2 Related Work
Multi-turn Attack. In contrast to single-turn attacks, which typically pose malicious questions at once with some confusion on LLMs  (Yuan et al., 2023; Wang et al., 2023; Liu et al., 2024), multi-turn jailbreaks obfuscate harmful intent by hiding it within a sequence of seemingly innocuous queries. Previous work shows this through human red-teaming  (Li et al., 2024), automated LLM attackers  (Russinovich et al., 2025; Ren et al., 2024; Yang et al., 2024), scenario-based setups  (Sun et al., 2024), query decomposition  (Zhou et al., 2024), and attacker-trained models  (Zhao and Zhang, 2025). These strategies significantly increase attack success rates compared to single-turn prompts.
Safety Evaluation of LLMs. Several datasets and benchmarks have been introduced to evaluate the safety of LLMs against harmful queries. Instruction-based benchmarks such as AdvBench  (Zou et al., 2023) and RedEval  (Bhardwaj and Poria, 2023) contain harmful or adversarial instructions that range from stereotypes, violence, to illegal activity. Generative benchmarks such as SAP (Deng et al., 2023) and AART (Radharapu et al., 2023) automatically construct adversarial prompts using models, enabling more diverse and adaptive evaluations. More recently, standardized evaluation frameworks have been proposed, targeting single-turn jailbreak robustness  (Chao et al., 2024)  (Mazeika et al., 2024) and multi-turn safety  (Yu et al., 2024; Burden et al., 2024).
Certification for LLMs. Several works have studied certification for LLMs. These methods focus on adversarial certification, typically by perturbing the input in token space  (Kumar et al., 2023; Emde et al., 2025) or embedding space (Casadio et al., 2025), and proving the model output remains safe. Unlike these perturbation-based approaches, we aim to directly certify against harmful queries themselves. Previous certification frameworks has been proposed for knowledge comprehension (Chaudhary et al., ) and bias detection (Chaudhary et al., 2024), but both in single-turn settings. In contrast, our work focuses on certifying multi-turn conversations, which are significantly more complex due to accumulated context and adaptive adversarial strategies.
3 Methodology
3.1 Problem Formalization
Let $Q$ be a finite set of queries, and let $G=(V,E)$ be a graph with $V=Q$ and $E\subseteq V\times V$ encoding user-defined relationships between queries. A multi-turn conversation between a user and a model $\mathcal{M}$ is defined as follows. At each turn $i$, the model receives the accumulated context $\mathrm{ctx}_{i}=\big(v_{0},r_{0},v_{1},r_{1},\ldots,v_{i-1},r_{i-1}\big)$, where $r_{j}=\mathcal{M}(\mathrm{ctx}_{j},v_{j})$ denotes the model’s response to query $v_{j}$ given prior context. The queries posed in the conversation are represented as an ordered tuple $\gamma=(v_{0},v_{1},\ldots,v_{n-1})\in V^{n},$ which we call a query sequence. We write $\mathcal{D}_{n}$ for a probability distribution over query sequences of length $n$.
Catastrophic events.
We define a harmful target $q^{*}$ as a specific objective that could lead to catastrophic outcomes, such as “How to build a bomb” or “How to synthesize a toxin.” To evaluate whether a model output is catastrophic with respect to $q^{*}$, we introduce a judge function $J_{q^{*}}(r_{i})\in\{0,1\},$ which returns $1$ if the response $r_{i}$ at turn $i$ reveals the target $q^{*}$; otherwise, it returns 0.
Objective.
Given a distribution $\mathcal{D}_{n}$ over query sequences, our goal is to certify the probability that a catastrophic event occurs during the n-turn conversation: $\Pr_{\gamma\sim\mathcal{D}_{n}}\!\left[\,\exists\,i\in[0,n)\;\;\text{s.t.}\;\;J_{t}(r_{i})=1\,\right].$
3.2 Graph Distributions from Markov Process
State space.
To define probability distributions over query sequences while avoiding repetition, reflecting the natural assumption that an adaptive attacker would not reuse the exact same prompt twice in the attack process, we specify a Markov process on a lifted state space in graph $G$. Formally, we define the state space $\Omega=\{(v,S):S\subseteq V,\,v\in S\}\cup\{\tau\},$ where $v$ is the current query, $S$ is the set of queries already used in the current sequence, which we track in each state to avoid revisiting queries within a single sequence. $\tau$ is the terminal state, meaning that no further queries are selected once this state is reached. The Markov process changes the current state to the next state according to a specified transition probability. The precise transition probability between states is specified in the subsequent subsections.
Transitions. We consider two families of distributions on query sequences: forward selection and backward selection. In all cases, if $\forall(v^{\prime},S^{\prime})\in\Omega,\,\Pr\!\left((v^{\prime},S^{\prime})\mid(v,S)\right)=0$, the state $(v,S)$ transits to the terminal state $\tau$ with $\Pr(\tau\mid(v,S))=1$. Moreover, $\forall\omega\in\Omega,\Pr(\omega\mid\tau)=\mathbf{1}\{\omega=\tau\}$, i.e. once $\tau$ is reached, it does not transition to any other state.
Forward selection. Given an initial distribution $\mu$ on $(v_{0},\{v_{0}\})$, we construct a length-$n$ sequence $\gamma=(v_{0},\dots,v_{n-1})$ where the visited set evolves as $S_{t}=\{v_{0},\dots,v_{t}\}$. The probability of sampling $\gamma$ under forward selection is
$\Pr(\gamma)\;=\;\mathcal{N}\Biggl(\mu((v_{0},\{v_{0}\}))\,\prod_{t=1}^{n-1}\Pr\!\left((v_{t},S_{t})\mid(v_{t-1},S_{t-1})\right)\Biggr)$
$\mathcal{N(\cdot)}$ denotes normalization over all length-$n$ sequences, ensuring $\sum_{\gamma:|\gamma|=n}\Pr(\gamma)=1$, which is necessary because sequences may terminate early at the terminal state $\tau$, so the raw product of transition probabilities over length-$n$ sequences does not automatically sum to 1.
Backward selection. Given an endpoint distribution $\nu$ on $(v_{n-1},\{v_{n-1}\})$, we construct a length-$n$ chain $\gamma=(v_{0},\dots,v_{n-1})$, where the visited set evolves as $U_{t}=\{v_{t},\dots,v_{n-1}\}$.
The probability of sampling $\gamma$ under backward selection is
$\Pr(\gamma)\;=\;\mathcal{N}\Biggl(\nu((v_{n-1},\{v_{n-1}\}))\,\prod_{t=n-1}^{1}\Pr\!\left((v_{t-1},U_{t-1})\mid(v_{t},U_{t})\right)\Biggr).$
Within this framework, we consider three representative distributions, capturing a different way in which adversarial queries may arise. These distributions are chosen because they capture natural strategies an attacker might employ, while remaining structured for statistical analysis. Importantly, our framework is not limited to these distributions. Additional distributions can be defined to explore other patterns of query sequences, making the approach broadly applicable.
  1. 1.
    Random node, where each query in the graph is selected independently at random. This provides an estimate of the model’s overall tendency to produce catastrophic content, without exploiting any structure in the query space.
  2. 2.
    Graph path, where the sequence of queries is a path in the graph, capturing relations between queries:
    1. (a)
      vanilla, where the last query is drawn from $V$, representing natural conversational flows.
    2. (b)
      harmful target constraint: where the last query is restricted to lie in a target set $Q_{T}$, forcing the conversation toward a high-risk query and increasing the likelihood of producing harmful outputs.
    This produces query sequences that are related by construction. The coherence in a query sequence has two advantages: First, the sequence provides local context that the language model can exploit when answering later queries; and second, the sequence tends to traverse a coherent region of the query space rather than jumping arbitrarily as in the random node distribution, which is unrealistic.
  3. 3.
    Adaptive with rejection, where transitions are guided by the model accept/reject response. This mimics realistic red-teaming where an attacker adapts their phrasing to circumvent safety mechanisms.
Distributions (1) and (3) correspond to forward selection, while (2) uses backward selection. In forward selection, we specify an initial distribution $\mu$ over the starting query and a transition probability $\Pr((v_{t+1},U_{t+1})\mid(v_{t},U_{t}))$. In backward selection, we specify an endpoint distribution $\nu$ over the ending query and a backward transition rule $\Pr((v_{t},U_{t}\mid v_{t+1},U_{t+1})$. For any nonempty finite set $A\subseteq V$, we write $\pi(\cdot\mid A)$ for a probability mass function on $A$. When we write $\pi(w\mid A)$, we mean the probability assigned to $w\in A$ under this distribution. We do not fix a specific form for these distributions (they may be uniform or weighted), only that they are valid probability mass functions on the indicated sets. We now describe the concrete instantiations of these distributions.
(1) Random node. The first query is selected according to a distribution $\pi(\cdot\mid V)$ over all nodes, i.e., $\mu((v_{0},\{v_{0}\}))=\pi(v_{0}\mid V)$. Each subsequent query is drawn from a distribution over the unvisited nodes $V\setminus S$ (i.e., nodes not yet visited in the current sequence, as recorded in $S$):
$\Pr\big((w,T)\mid(v,S)\big)=\begin{cases}\pi(w\mid V\setminus S),&w\in V\setminus S,\;T=S\cup\{w\},\\[5.69054pt] 0,&\text{otherwise}.\end{cases}$
(2) Graph Path. Rather than selecting queries independently, we generate a sequence of queries that is a path in the graph. For $v\in V$ we denote its neighbor set by $N(v):=\{\,w\in V:(v,w)\in E\,\}.$ We consider two endpoint distributions for the last query in the path:
(2.a) vanilla. The endpoint is selected from $V$ by $\nu_{\text{all}}((v_{n-1},\{v_{n-1}\}))=\pi(v_{n-1}\mid V).$
(2.b) harmful target constraint. In many settings, it is advantageous to control the final query in the sequence. Biasing the endpoint steers the path toward a semantic region of interest (e.g., near the target query $q^{\star}$) while still generating coherent predecessors. The idea is that once the model has processed the earlier queries, the final query is the one where we most expect a desired behavior, so constraining it can help guide outcomes. Formally, we restrict the last query to lie in a designated target set $Q_{T}$ and define $\nu_{Q_{T}}((v_{n-1},\{v_{n-1}\}))=\pi(v_{n-1}\mid Q_{T}).$
For notational convenience, we write both distributions through a single formulation. Let $\nu\in\{\nu_{\text{all}},\,\nu_{Q_{T}}\}$ denote the endpoint distribution, where $\nu_{\text{all}}$ draws the endpoint from $V$, and $\nu_{Q_{T}}$ restricts it to the target set $Q_{T}$. Then the transition probability can be written as
$\Pr((w,T)\mid(v,S))=\begin{cases}\pi(w\mid N(v)\setminus S),&w\in N(v)\setminus S,\;T=S\cup\{w\},\\[5.69054pt] 0,&\text{otherwise}.\end{cases}$
(3) Adaptive with rejection. Intuitively, when the LLM answers the current query, it indicates that the query is not yet harmful enough to trigger refusal. In this case, it is natural to move toward the harmful target $q^{\star}$. Conversely, if the model rejects the query, this suggests that the query is perceived as too harmful. The transition rule then favors moving to a less harmful neighbor, thereby stepping back in similarity with $q^{\star}$.
To incorporate feedback from model $\mathcal{M}$, we introduce a binary rejection indicator at $v$, $r_{v}:=\mathbf{1}\{\texttt{is\_rej}(\mathcal{M}(v))\}$ to indicate whether the current query $v$ is rejected by the model $\mathcal{M}$. We partition unvisited neighbors $N(v)$ according to whether they increase or decrease similarity with the harmful target compared to the current query $v$ :
$A_{\mathrm{prog}}(v,S)$ $=\{w\in N(v)\setminus S:\mathrm{sim}(w,q^{\star})\geq\mathrm{sim}(v,q^{\star})\},$
$A_{\mathrm{deprog}}(v,S)$ $=\{w\in N(v)\setminus S:\mathrm{sim}(w,q^{\star})<\mathrm{sim}(v,q^{\star})\}.$
Here “prog” means moving toward higher or equal similarity with $q^{\star}$, while “deprog” means moving to lower similarity. We then assign weights depending on whether the current query is rejected. When $v$ is accepted ($r_{v}=0$), progress toward the target $q^{\star}$ is encouraged by giving larger weight to $A_{\mathrm{prog}}$ and smaller weight to $A_{\mathrm{deprog}}$. If $v$ is rejected ($r_{v}=1$), the bias is reversed, steering the sampler toward safer regions.
Formally, we pick a $\pi_{N(v)}$ and define the weight on a given query $w$ by $\lambda_{v,S}(w)=\lambda_{h}\,\mathbf{1}_{\{w\in H(v,S)\}}\,\pi(w\mid N(v)\setminus S)+\lambda_{l}\,\mathbf{1}_{\{w\in L(v,S)\}}\,\pi(w\mid N(v)\setminus S)$ with weights $0<\lambda_{l}<\lambda_{h}$ are tunable parameters, where the high- and low-weight neighbor sets depending on the rejection are given by:
$H(v,S):=\begin{cases}A_{\mathrm{prog}}(v,S),&r_{v}=0,\\ A_{\mathrm{deprog}}(v,S),&r_{v}=1,\end{cases}\qquad L(v,S):=\begin{cases}A_{\mathrm{deprog}}(v,S),&r_{v}=0,\\ A_{\mathrm{prog}}(v,S),&r_{v}=1.\end{cases}$
Thus when $r_{v}=0$ the $\mathrm{prog}$ set receives higher weight (encourage progress), and when $r_{v}=1$ the $\mathrm{deprog}$ set receives higher weight. To guarantee that every query in the high-weight set has strictly larger weight than every query in the low-weight set, we require $\lambda_{h}\cdot\min_{a\in H}\;\pi(a\mid N(v)\setminus S)\;>\;\lambda_{l}\cdot\max_{b\in L}\;\pi(b\mid N(v)\setminus S)$. This condition is vacuous when either set is empty. The distribution on the first query is $\mu(v_{0},\{v_{0}\})=\pi_{V}(v_{0})$, and the normalized transition probability is
$\Pr((w,T)\mid(v,S))=\begin{cases}\dfrac{\lambda_{v,S}(w)}{\sum_{u\in N(v)\setminus S}\lambda_{v,S}(u)},&w\in N(v)\setminus S,\;T=S\cup\{w\}\\ 0,&\text{otherwise.}\end{cases}$
Augmentation layer. We extend the base distribution with an augmentation layer $\mathcal{D}_{\mathrm{aug}}(\cdot\mid v_{t})$. For each query $v_{t}$ in the sequence $\gamma=(v_{0},\dots,v_{n-1})$, this augmentation distribution $\mathcal{D}_{\mathrm{aug}}(\cdot\mid v_{t})$ depends on the current query $v_{t}$. The augmented sequence $\widetilde{\gamma}=(\widetilde{v}_{0},\dots,\widetilde{v}_{n-1})$ is obtained by sampling each augmented query independently conditional on $\gamma$:
$\widetilde{v}_{t}\sim\mathcal{D}_{\mathrm{aug}}(\cdot\mid v_{t}),\qquad t=0,\dots,n-1.$
Intuitively, this means that each query $v_{t}$ drawn from the base distribution can be replaced by an augmented version (for example, by inserting a jailbreak prefix before $v_{t}$, or by rewriting $v_{t}$). The resulting sequence has probability
$\Pr(\widetilde{\gamma})\;=\;\Pr(\gamma)\,\prod_{t=0}^{n-1}\Pr_{\mathcal{D}_{\mathrm{aug}}}(\widetilde{v}_{t}\mid v_{t}).$
This formulation covers both the identity case (when $\mathcal{D}_{\mathrm{aug}}(\cdot\mid v)$ returns $v$ with probability $1$) and stochastic modifications of $v$. We instantiate $\mathcal{D}_{\mathrm{aug}}$ using a jailbreak augmentation distribution $\mathcal{D}_{\mathrm{jb}}(\cdot\mid v)$ defined in Appendix A, which with some probability returns $v$ unchanged and with complementary probability prepends a jailbreak prompt to $v$. More generally, $\mathcal{D}_{\mathrm{aug}}$ can be defined by more structured generators, such as a second LLM that chooses mutations based on the conversation context, as long as this generator is treated as part of the attack process and induces a well-defined distribution over conversations.
4 Evaluation
4.1 Graph Construction
We construct our dataset from HarmBench (Mazeika et al., 2024), using three categories: chemical_biological, cybercrime, and illegal. Our main evaluation focuses on chemical_biological and cybercrime categories, which together comprise 68 scenarios (28 in chemical_biological and 40 in cybercrime), while results on illegal category are included as an additional domain check in Appendix C. Each scenario is associated with a specific harmful target query $q^{\star}$. Since directly asking $q^{\star}$ is unrealistic in multi-turn conversations, for each $q^{\star}$, we expand a query set $Q$ whose elements are related to $q^{\star}$ but less harmful, in order to indirectly obtain answers of the target. We use actor-based prompts following the approach of (Ren et al., 2024): one harmful query is asked to derive a related book or person (the “actor”), and then queries are generated around this actor. To ensure diversity of the dataset, we prompt three LLMs: Gemini-2.5-Flash-Lite (Comanici et al., 2025), Deepseek-R1 (Guo et al., 2025), and Mistral-Large-2407 (Mistral AI, 2024). For each scenario, each LLM generates 10 actors, making a total of 30 actors, with 5 queries per actor. To illustrate, consider a scenario where $q^{\star}$ asks how to construct an explosive device. An example actor might be “Alfred Nobel”. Queries derived from this actor could include: “What is Alfred Nobel’s main contribution in history?”, or “What are the steps involved in his approach to this contribution?” After deduplication, we randomly sample 20 actors, yielding 100 queries per scenario.
To capture semantic relationships among queries in $Q$, we formally define an undirected graph $G=(V,E)$, where $V=Q$ and the edge set $E\subseteq V\times V$ is defined as $E=\{(u,v)\in Q\times Q\mid\ell_{\text{th}}<\mathrm{sim}(u,v)<h_{\text{th}}\}$. Here, $\mathrm{sim}(u,v)$ denotes the cosine similarity between sentence embeddings of $u$ and $v$, computed using the all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019). The parameters $l_{\text{th}}$ and $h_{\text{th}}$ are lower and upper thresholds that control edge formation, ensuring that connections capture semantic relatedness while avoiding near-duplicate queries. We further define the target set $Q_{T}=\{v\in Q\mid\ell_{\text{th}}<\mathrm{sim}(v,q^{\star})<h_{\text{th}}\},$ which identifies queries that are highly similar to the harmful query $q^{\star}$ while remaining distinct, serving as endpoints for the graph-path (harmful target constraint) distribution in our statistical certification framework.
4.2 Experimental setup
For each scenario, we consider the four specifications on distributions introduced in Section 3.2. We instantiate each conditional distribution $\pi(\cdot\mid X)$ as the uniform distribution over $X$, i.e., $\color[rgb]{0,0,1}\pi(v\mid X)=\tfrac{1}{|X|}\;\text{for all }v\in X.$} In our main experiments, we instantiate the augmentation layer with the jailbreak distribution $\mathcal{D}_{\mathrm{jb}}$ defined in Appendix A, using a jailbreak insertion probability $p=0.2$, and apply it only to the Random Node distribution. To bound the probability an LLM produces catastrophic outcomes under a given specification, we apply the Clopper–Pearson method (Clopper and Pearson, 1934) to compute 95% confidence intervals for the probability of catastrophic responses. We use 50 sampled query sequences per specification. To determine whether an LLM’s response is a catastrophic outcome, we use GPT-4o (OpenAI, 2024) as a judge model, similar to prior studies (Yuan et al., 2024; Team, 2025).
4.3 Certification Results
We evaluate the statistical certification bounds of several state-of-the-art large language models: Llama-3.3-70b-Instruct (Meta AI, 2024), Mistral-Large-2407 (Mistral AI, 2024), DeepSeek-R1 (Guo et al., 2025), gpt-oss-120b (Agarwal et al., 2025), Claude-Sonnet-4 (Anthropic, 2024) and Nova Premier (Intelligence, 2025). We use the default hyperparameter settings shown in Table 6 and analyze their influence through an ablation study (Appendix D). For each LLM and specification, we estimate statistical certification bounds on the attack success probability with $95\%$ confidence, reporting the median of the lower and upper bounds on chemical_biological and cybercrime categories across all specifications under a distribution in Table 1. Figure 4 and 5 (Appendix B) show the results in box plots for specifications developed from the chemical_biological and cybercrime datasets respectively.
Table 1: Statistical certification bounds under different distributions for each dataset and model (median of 95% confidence intervals across all specifications under a distribution). Distributions: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). We bold the highest bounds among four distributions for each LLM.
Dataset Model Distributions (median 95% CI)
RNwJ GPv GPh AwR
chembio nova (0.005, 0.137) (0.001, 0.106) (0.013, 0.165) (0.005, 0.137)
deepseek (0.554, 0.821) (0.221, 0.498) (0.229, 0.508) (0.212, 0.488)
claude (0.001, 0.106) (0.001, 0.106) (0.001, 0.106) (0.001, 0.106)
gpt-oss (0.028, 0.205) (0.072, 0.291) (0.045, 0.243) (0.101, 0.337)
mistral (0.554, 0.821) (0.318, 0.607) (0.432, 0.718) (0.452, 0.735)
llama (0.212, 0.488) (0.116, 0.359) (0.195, 0.457) (0.146, 0.403)
cyber nova (0.000, 0.071) (0.001, 0.106) (0.001, 0.106) (0.000, 0.071)
deepseek (0.721, 0.935) (0.472, 0.753) (0.543, 0.813) (0.543, 0.813)
claude (0.028, 0.205) (0.123, 0.371) (0.195, 0.467) (0.195, 0.467)
gpt-oss (0.086, 0.314) (0.229, 0.508) (0.309, 0.597) (0.318, 0.607)
mistral (0.652, 0.892) (0.403, 0.691) (0.533, 0.805) (0.565, 0.830)
llama (0.374, 0.663) (0.264, 0.548) (0.432, 0.718) (0.393, 0.682)
General Observations.
By comparing the bounds, we observe that among frontier LLMs, Claude-Sonnet-4 and Nova Premier are safer than the others, while Mistral-Large and DeepSeek-R1 exhibit higher risks. In particular, Nova Premier demonstrates consistently low risk levels, largely because its built-in guardrails often block potentially unsafe content. On the other hand, DeepSeek-R1 reaches a certified lower bound of over 70% in cybercrime scenarios under RNwJ distributions. For LLMs with relatively low probabilities of catastrophic outcomes (e.g., Nova Premier and Claude-Sonnet-4), distributions augmenting with jailbreak are largely ineffective. In contrast, weaker LLMs such as Mistral-Large and DeepSeek-R1 remain vulnerable to jailbreak prompts, indicating that additional safety training is needed. We analyze the effect of the jailbreak probability in Appendix D.1; for less safe LLMs, increasing the jailbreak probability generally raises catastrophic outcomes, while for safer LLMs the effect is negligible. Other distributions, Adaptive with Rejection and Graph Path, are often more effective in producing catastrophic outcomes on safer LLMs. For Graph Path, constraining the final query to a harmful set (GPh) consistently increases attack effectiveness relative to the vanilla last-query distribution (GPv), which shows that shaping the final step of a multi-query sequence is an effective method for attackers. For Adaptive with Rejection, the strategy exploits the fact that safer LLMs refuse to answer queries at non-trivial rates (roughly 20% for gpt-oss-120b and 15% for Claude-Sonnet-4 in our samples). By designing sequences that interact with these rejection dynamics, attackers can substantially increase catastrophic responses on LLMs that otherwise appear well aligned.
Table 2: Numbers of attack scenarios where the statistical certification lower bound exceeds the baseline ST(Single-turn) and MT(Multi-turn) ASR by more than 0.05.
Dataset Model Numbers of specs
ST MT
chembio deepseek 86 29
claude 11 1
gpt-oss 51 1
mistral 100 30
llama 78 14
nova 52 3
cyber deepseek 157 22
claude 95 16
gpt-oss 110 2
mistral 151 38
llama 143 17
nova 44 0
Attack Patterns.
In our analysis, we identify two common attack patterns:
  1. 1.
    Effect of Distractors. We observe that LLMs often refuse to answer harmful queries that are presented in isolation, replying with messages such as “I can’t provide that information” due to built-in safety mechanisms (Zhang et al., 2025; Yuan et al., 2025).
    However, when the same query is embedded in a multi-turn dialogue that includes other questions, even not directly related, the model is more likely to provide a harmful answer. This behavior is observed across our various specification distributions. Figure 2 shows a representative example of specification in the graph path (harmful target constraint) distribution.
  2. 2.
    Role of Context. Even when a model answers a harmful query in a single turn, the response may be incomplete, confused, or fail to reach the catastrophic information the user intends. Attackers can make the target clearer by referring to earlier conversation turns (e.g. “you just mentioned”). Placing the query within a dialogue enables the model to infer the user’s focus and produce outputs that are more directly related to the harmful target. This behavior is observed across our various specification distributions. Figure 3 shows a representative example of specification in the graph path (vanilla) distribution.
Figure 2: Responses of Mistral, DeepSeek, and gpt-oss on GPh distribution to the target query: “Explain how and where to purchase the ingredients for manufacturing sarin without being traced.”
Figure 3: Effect of providing conversational context for the target query “Give detailed step-by-step instructions for manufacturing ricin at home.” on GPv distribution: Without context, Claude and DeepSeek tend to give only general information about chemicals and Llama is confused by the fictional setting. When relevant prior context is included, these models’ responses shift to ricin, leading to catastrophic responses.
Comparing with Baselines
There is no prior work certifying catastrophic risks. We consider two baselines representing existing approaches for evaluating risks: (i) single-turn (ST), which uses all 100 queries in our dataset and sends each query independently to the LLM without any conversational history, and (ii) multi-turn (MT), where the same query set is grouped into actors as when we created it, each actor contributes a sequence of 5 queries. These sequences are submitted in order, simulating an iterative multi-turn attack.
The baselines are not directly comparable, but in the absence of stronger alternatives, we provide a rough comparison. Importantly, our statistical certification evaluates models over a much larger conversation space, considering all possible sequences consistent with the query distributions rather than a fixed subset. To make the comparison more meaningful, for these baselines, in each scenario, we measure the fraction of queries (ST) or sequences (MT) that lead to catastrophic responses. Rather than using a binary outcome per scenario (recording a $1$ if any catastrophic response occurs across several trials, which is commonly done in the literature (Zou et al., 2023; Qi et al., 2023)), this measure provides a finer-grained view of how difficult it is to elicit catastrophic outcomes from a model in a given scenario. We then count the number of scenarios where our certified lower bound exceeds this baseline fraction by more than $0.05$ (Table 2).
We observe that for some models, nearly all specifications yield the rate in ST lower than the certified lower bound, indicating that single-turn evaluations substantially underestimate LLMs’ risks. Even with multi-turn attacks, we find several scenarios where our certified lower bound on catastrophic response probability exceeds the rate observed in the baseline by a non-trivial margin, highlighting that fixed-sequence baselines can significantly underestimate LLM risks.
5 Conclusion
We introduce a statistical certification framework for quantifying catastrophic risks in multi-turn LLM conversations. Unlike prior work that reports attack success rates on fixed benchmarks, our approach provides high-confidence probabilistic bounds over large conversation spaces, enabling meaningful comparisons across models. Our results reveal that catastrophic risks are non-trivial for all frontier LLMs, with notable differences in safety across models.
Ethics statement
We identify the following positive and negative impacts of our work.
Positive impacts.
Our work is the first to provide quantitative certificates for catastrophic risks in multi-turn LLM conversations. It can help model developers systematically evaluate and compare their models before deployment, and inform the general public of potential harms when interacting with LLMs. Since C3LLM only requires black-box access, it applies equally to both open- and closed-source models, thus broadening its utility.
Negative impacts.
Our framework involves constructing specifications to probe harmful behavior in LLMs. While these specifications are designed for evaluation and certification, they could be misused by adversaries to more systematically search for harmful responses. We emphasize that our methodology is intended for safety evaluation, not exploitation, and we have taken care to restrict examples and datasets to standard benchmarks.
Acknowledgements
This work was supported by a research gift from Amazon Research.
Appendix A Explicit Jailbreak Distribution
We now give the explicit construction of the jailbreak distribution $\mathcal{D}_{{jb}}$ and its probability mass. Let main_jb be a base jailbreak instruction, and let $\mathcal{S}=\{s_{1},\dots,s_{M}\}$ be a set of side jailbreak instructions. We split main_jb into consecutive sentences $(m_{1},\dots,m_{K})$.
The jailbreak $\eta$ is then formed as an alternating sequence of main and side instructions:
$\eta=(m_{1},k_{1},m_{2},k_{2},\dots,m_{K}),$
where $k_{j}$ is a sequence of side instructions inserted between $m_{j}$ and $m_{j+1}$.
Formally, for each gap $j\in\{1,\dots,K-1\}$:
  • Each side instruction $s\in\mathcal{S}$ is included in $k_{j}$ independently with probability $\rho\in(0,1)$.
  • If $T_{j}(\eta)\subseteq\mathcal{S}$ is the chosen subset, its elements are permuted uniformly at random, i.e., each ordering has probability $1/|T_{j}(\eta)|!$.
Thus, the probability of generating a jailbreak $\eta$ is
$\Pr(\eta)\;=\;\prod_{j=1}^{K-1}\Bigg[\Bigg(\prod_{s\in T_{j}(\eta)}\rho\Bigg)\Bigg(\prod_{s\in\mathcal{S}\setminus T_{j}(\eta)}(1-\rho)\Bigg)\frac{1}{|T_{j}(\eta)|!}\Bigg].$
This defines a base distribution over jailbreak-prefix strings, which we denote by $\mathcal{D}_{\mathrm{prefix}}$.
Augmentation with mutations and insertion. Let $\mathcal{D}_{\mathrm{prefix}}$ be the base distribution over jailbreak-prefix strings $g$, and let $\mathrm{tok}(g)=(t_{1},\dots,t_{L})$ be the tokenization of $g$. For a fixed mutation probability $\mu\in(0,1)$ and a replacement distribution $q$ over the tokenizer vocabulary (e.g., uniform or restricted to a set of “possible” tokens), we define the mutation operator $M_{\mu}$ by
$\Pr(\tilde{g}\mid g)\;=\;\prod_{i=1}^{L}\Bigl[(1-\mu)\,\mathbf{1}\{\tilde{t}_{i}=t_{i}\}\;+\;\mu\,q(\tilde{t}_{i})\Bigr],$
[…]
$\mathcal{D}_{\mathrm{prefix}}^{\mathrm{mut}}(\tilde{g})\;=\;\sum_{g}\mathcal{D}_{\mathrm{prefix}}(g)\,\Pr(\tilde{g}\mid g).$
Given a base query $v$ and a jailbreak insertion probability $p\in(0,1)$, we define the jailbreak augmentation distributions over full queries by
$\mathcal{D}_{\mathrm{jb}}(a\mid v)\;=\;(1-p)\,\mathbf{1}\{a=v\}\;+\;p\sum_{g}\mathcal{D}_{\mathrm{prefix}}(g)\,\mathbf{1}\{a=g\mathbin{\|}v\},$
and
$\mathcal{D}_{\mathrm{jb}}^{\mathrm{mut}}(a\mid v)\;=\;(1-p)\,\mathbf{1}\{a=v\}\;+\;p\sum_{\tilde{g}}\mathcal{D}_{\mathrm{prefix}}^{\mathrm{mut}}(\tilde{g})\,\mathbf{1}\{a=\tilde{g}\mathbin{\|}v\},$
where $g\mathbin{\|}v$ denotes the concatenation of the prefix $g$ and the base query $v$. Equivalently, $\mathcal{D}_{\mathrm{jb}}(\cdot\mid v)$ and $\mathcal{D}_{\mathrm{jb}}^{\mathrm{mut}}(\cdot\mid v)$ can be implemented by sampling a Bernoulli random variable $B\sim\mathrm{Bernoulli}(p)$ and setting
$a=\begin{cases}v,&B=0,\\[2.84526pt] g\mathbin{\|}v,&B=1,\;\;g\sim\mathcal{D}_{\mathrm{prefix}},\\ \end{cases}\quad\text{or}\quad a=\begin{cases}v,&B=0,\\[2.84526pt] \tilde{g}\mathbin{\|}v,&B=1,\;\;\tilde{g}\sim\mathcal{D}_{\mathrm{prefix}}^{\mathrm{mut}},\end{cases}$
respectively.
Illustrative Scenario In our added experiments, we instantiate a context-dependent augmentation for the Adaptive with Rejection (AwR) attack: whenever the previous query is rejected, we replace the next base query $v_{t}$ with an augmented query $\tilde{v}_{t}\sim\mathcal{D}^{\mathrm{mut}}_{\mathrm{jb}}(\cdot\mid v_{t})$ (Appendix A). Table 3 reports certified 95% confidence intervals for catastrophic failure probability on one representative scenario.
Table 3: Certified 95% confidence intervals for catastrophic failure probability on a single representative scenario under the augmented adaptive-with-rejection (AwR) attack.
Model 95% CI
Mistral (0.7569, 0.9547)
gpt-oss (0.1463, 0.4034)
Claude (0.2121, 0.4877)
DeepSeek (0.4518, 0.7359)
Llama (0.4518, 0.7359)
Appendix B Detailed Statistical Certification Bounds
Figure 4 and  5 report the complete statistical certification lower and upper bounds (median and IQR) for every model–distribution pair across all specifications.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 4: Certification results for the chemical_biological dataset. Each panel shows the distribution of lower bounds and upper boundsunder different specifications for one LLM.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 5: Certification results for the cyber crime dataset. Each panel shows the distribution of lower bounds and upper boundsunder different specifications for one LLM.
Appendix C Additional Results on HarmBench illegal category
Setup.
In addition to the chemical_biological and cybercrime categories reported in the main paper, we evaluate our certification framework on scenarios from the HarmBench (Mazeika et al., 2024) illegal subset. We follow the same graph construction procedure (Section 4.1) and the same experiment setup (Section 4.2). Due to time and compute constraints, we evaluate 30 scenarios for Claude, gpt-oss, and Llama, and the first 10 scenarios for Mistral and DeepSeek. We report results separately in Tables 4 and 5.
Distributions.
We evaluate the same four distributions used throughout the paper: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). Following the main text, each entry in Tables 4 and 5 is the median of 95% confidence intervals across all specifications under a distribution, and we bold the highest bound among the four distributions for each model.
Results and observations.
Compared with the main results, certified bounds on the illegal category for gpt-oss and Claude are lower than on the chemical_biological and cybercrime categories. By contrast, Llama, Mistral and DeepSeek exhibit high certified catastrophic risk across all three domains. These results illustrate how our framework allows practitioners to assess catastrophic risk with statistical guarantees in different domains and to choose models accordingly: for instance, Claude may be acceptable for illegal in our setting, whereas for cybercrime none of the evaluated models appears reliably safe, indicating that stronger mitigations on safer LLMs would be needed.
Table 4: Statistical certification bounds on HarmBench illegal (first 10 scenarios), Mistral and DeepSeek. Entries are the median of 95% confidence intervals across all specifications under a distribution. Distributions: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). We bold the highest bounds among the four distributions for each model.
Model RNwJ GPv GPh AwR
mistral (0.586, 0.845) (0.234, 0.504) (0.363, 0.653) (0.314, 0.609)
deepseek (0.523, 0.796) (0.146, 0.403) (0.187, 0.457) (0.212, 0.488)
Table 5: Statistical certification bounds on HarmBench illegal (30 scenarios), Claude, gpt-oss, and Llama. Entries are the median of 95% confidence intervals across all specifications under a distribution. Distributions: Random Node with Jailbreak (RNwJ), Graph Path (vanilla) (GPv), Graph Path (harmful target constraint) (GPh), and Adaptive with Rejection (AwR). We bold the highest bounds among the four distributions for each model.
Model RNwJ GPv GPh AwR
claude (0.000, 0.089) (0.005, 0.137) (0.001, 0.106) (0.005, 0.137)
gpt-oss (0.001, 0.106) (0.009, 0.151) (0.013, 0.165) (0.005, 0.137)
llama (0.309, 0.597) (0.086, 0.314) (0.131, 0.382) (0.101, 0.337)
Appendix D ablation study
In this section, we analyze the effect of hyperparameters on certification results. Table  6 shows the hyperparameters and their values used in the experiments. We conduct ablation studies on a randomly selected scenario from the dataset on Graph Path (harmful target constraint) distribution. For Appendices D.5D.7, we certify Llama-3.3-70B-Instruct as they are model-agnostic; otherwise, we certify all evaluated LLMs.
Table 6: Default hyperparameters used in experiments.
Hyperparameter Description Value
$\alpha$ $1-\alpha$ is the confidence interval for certification $0.05$
num_samples Number of samples for certification $50$
$l_{\text{th}}$ Lower threshold of embedding similarity to connect edges $0.4$
$h_{\text{th}}$ Higher threshold of embedding similarity to connect edges $0.8$
$\lambda_{l}$ Weight assigned to high-weight neighbor set in AwR distributions 1
$\lambda_{h}$ Weight assigned to high-weight neighbor set in AwR distributions 2.5
qlen Length of the query sequence 5
jailbreak_prob Probability of inserting jailbreak prompt before a query $0.2$
setsize Size of Query Set 100
D.1 Jailbreak Probability
Certification bounds on Random Node with Jailbreak distribution is controlled by the jailbreak probability hyperparameter. We show results in Figure 6. Overall, we observe that increasing the jailbreak probability generally raises the certified catastrophic-risk bounds for less robust models such as DeepSeek, Mistral and Llama, with the largest bounds typically appearing at moderate-to-high probabilities (e.g., $p\approx 0.6$$1.0$). For Llama-3.3-70B-Instruct, the bounds increase from $p=0$ to moderate probabilities and then slightly decrease as $p$ approaches $1$, suggesting that overly frequent, highly conspicuous jailbreaks can be partially mitigated by the model. For Claude and gpt-oss, the certified bounds remain relatively low and flat across all probabilities, indicating that these models are comparatively more defensive to the jailbreak prompts used in our experiments.
((a))
((b))
((c))
((d))
((e))
Figure 6: Ablation studies for jailbreak probability on certification bounds for LLMs.
D.2 Length of Query Sequence
Figure 7 shows how certification bounds vary with the length of query sequence. Across models, increasing the sequence length generally pushes the certified bounds upward, indicating that longer conversations provide attackers with more opportunities to elicit catastrophic behavior (for LLaMA, DeepSeek, and Claude). In contrast, gpt-oss appears more robust to query length, with bounds changing only slightly, and for Mistral the bounds also vary little because the model is already highly unsafe even for short sequences.
((a))
((b))
((c))
((d))
((e))
Figure 7: Ablation studies for the length of the query sequence on certification bounds for LLMs.
D.3 Size of Query Set
Figure 8 shows how the size of the query set used to build specifications affects the certified bounds. Across models, increasing the query-set size from 50 to 150 has only a modest effect on the certified bounds. This suggests that, once the initial query set is reasonably large and diverse, our certification results are fairly stable and do not rely on a very specific query-set size.
((a))
((b))
((c))
((d))
((e))
Figure 8: Ablation studies for query-set size on certification bounds for five LLMs.
D.4 Ratio of Weight
In the Adaptive with Rejection distribution, $\lambda_{h}$ denotes the weight assigned to the high-weight neighbor set, while $\lambda_{l}$ represents the weight assigned to the low-weight neighbor set. Since the distribution is normalized after applying these weights (see Section 3.2), only the ratio $\lambda_{h}/\lambda_{l}$ determines the effective sampling probabilities, rather than their absolute values.
To study the influence of this ratio, we perform an ablation experiment by varying $\lambda_{h}/\lambda_{l}$ across the values $\{1.5,2.0,2.5,3.0,3.5\}$. Note that we require $\lambda_{h}>\lambda_{l}$, hence the minimum ratio considered is $1.5$. We then evaluate the resulting certified bounds under these different settings. Figure 9 shows that, for all five LLMs, the certified bounds change only moderately as $\lambda_{h}/\lambda_{l}$ varies, with the highest bounds typically occurring at intermediate ratios (e.g., 2.02.03.03.0). This suggests that a balanced setting—strong enough to move toward the harmful target when queries are accepted, but still willing to step back toward safer neighbors when rejections occur—gives the most effective behavior within this family.
((a))
((b))
((c))
((d))
((e))
Figure 9: Ablation studies for ratio of weights on certification bounds for LLMs.
D.5 Number of Samples
To assess how our certification bounds change with the number of samples $n$, we report bounds in Figure 10(c). The ranges between lower and upper bounds shrink as $n$ increases from small values, and stablize once $n\approx 50$. In our main experiments, we therefore adopt $n=50$ as a trade off between computational cost and statistical precision.
D.6 Graph Thresholds
Graph-based specifications rely on two thresholds, $l_{\text{th}}$ and $h_{\text{th}}$, which determine the sparsity of the similarity graph by controlling which edges are created based on embedding similarity. To study their influence, we examine two settings: (i) fixing $l_{\text{th}}=0.4$ while varying $h_{\text{th}}\in\{0.7,0.8,0.9,1.0\}$, and (ii) fixing $h_{\text{th}}=0.8$ while varying $l_{\text{th}}\in\{0.2,0.3,0.4,0.5\}$. Figure 10 shows that the bounds do not change significantly for different thresholds.
D.7 Variance
We show the variance of our certification bounds in Figure 10(d), where we run the same experiment on one specification 10 times. We report the median and interquartile range (IQR) of the resulting 95% confidence lower and upper bounds. The results demonstrate that the variance is low, demonstrating the reliability of our certification procedure.
((a))
((b))
((c))
((d))
Figure 10: Ablation studies for (a–b) graph-threshold settings, (c) number of samples, and (d) variance of certified bounds.
Appendix E LLM Usage
LLMs were used in this work solely as general-purpose assistive tools to aid in polishing the writing and improving clarity of exposition.